The microchip production times at our different suppliers are distributed N(44 , 10) (hrs.). To study this distribution, we use Minitab to randomly select a sample of 100 numbers from this distribution. Use this information to answer Questions 6 & 7.

6. We expect the standard deviation of these 100 numbers to be:

A. 10 B. 1.0 C. 0.10

7. For purposes of our study, we define an outlier to be any number in the sample less than 24 or greater than 64. Using the empirical rule, approximately how many outliers do we expect in this sample?

A. 5 B. 10 C. 16 D. 0

What's the solution step of Q7?
Answer #1

6)\ We expect the standard deviation of these 100 numbers to be =10 ; option A

7)as % of values within 2 standard deviation from mean (24 to 64)=95.0%

therefrore % of outliers =100-95 =5%

approximately number of outliers =100*5%=5

option A